40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14471     Entry time: Wed Feb 27 21:34:21 2019
Author: gautam 
Type: Update 
Category: General 
Subject: Suspension diagnosis 

In my effort to understand what's going on with the suspensions, I've kicked all the suspensions and shutdown the watchdogs at 1235366912. PSL shutter is closed to avoid trying to lock to the swinging cavity. The primary aims are

  1. To see how much the resonant peaks have shifted w.r.t. the database, if at all - I claim that the ETMY resonances have shifted by a large amount and also has lost one of the resonant peaks.
  2. To check the status of the existing diagonalization.

All the tests I have done so far (looking at free swinging data, resonant frequencies in the Oplev error signals etc) seem to suggest that the problem is mechanical rather than electrical. I'll do a quick check of the OSEM PD whitening unit in 1Y4 to be sure.But the fact that the same three peaks appear in the OSEM and Oplev spectra suggests to me that the problem is not electrical.

Watchdogs restored at 10 AM PST

ELOG V3.1.3-